OpenStudy (anonymous):

Evaluate the function rule to find the range of each function for the domain {-3, 0, 5}. f(x) = 4x − 9 A. {-21, -9, 11} B. {-21, 9, 29} C. {-9, 3, 29} D. {3, 9, 21}

OpenStudy (anonymous):

someone help pl

OpenStudy (anonymous):

please!!!!!!

OpenStudy (anonymous):

x =1/4y + 9/4

OpenStudy (anonymous):

This is how i get the answer Let's solve for x. y=4x−9 Step 1: Flip the equation. 4x−9=y Step 2: Add 9 to both sides. 4x−9+9=y+9 4x=y+9 Step 3: Divide both sides by 4. 4x 4 = y+9 4 x= 1 4 y+ 9 4

OpenStudy (anonymous):

i think its B?

OpenStudy (anonymous):

r u there

OpenStudy (anonymous):

HELP ON THE VERGE OF CRYING PLEASE HELP!!!!

OpenStudy (saifoo.khan):

Do you know what are domains and range?

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I was helping someone else

OpenStudy (anonymous):

not really no?

OpenStudy (anonymous):

@saifoo.khan do you have this one?

OpenStudy (saifoo.khan):

Nope. @Thehulk49 you may continue.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

do one if u know this answer? please

OpenStudy (anonymous):

Answer: x =1/41 + 9/4

OpenStudy (anonymous):

that not a choice

OpenStudy (anonymous):

A. {-21, -9, 11} B. {-21, 9, 29} C. {-9, 3, 29} D. {3, 9, 21}

OpenStudy (anonymous):

I wan to say c but i am not sure

OpenStudy (anonymous):

do u have another guess

OpenStudy (anonymous):

can u help with another?

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

Evaluate the function rule to find the range of each function for the domain {-3, 0, 5}. y = -2x + 6 A. {-6, -4, 0} B. {-4, 0, 6} C. {-4, 6, 12} D. {6, 12, 16}